# Large Language Model Bias Mitigation from the Perspective of Knowledge Editing

### v1

@@ 1 Introduction
Pre-trained Large Language Models (LLMs) have demonstrated exceptional performance on many tasks (Devlin et al., 2018; Floridi & Chiriatti, 2020; Brown et al., 2020). However, the encoded social stereotypes and human-like biases inevitably cause undesired behaviors when deploying LLMs in practice (Zhao et al., 2019; Navigli et al., 2023; Sheng et al., 2021). Existing approaches to mitigate biases in LLMs are mainly categorized into: (1) Fine-tuning (Zmigrod et al., 2019; Webster et al., 2020; He et al., 2022; Liang et al., 2020; Lauscher et al., 2021), which includes techniques such as re-balanced corpus pre-training, contrastive learning, projection methods, and efficient parameter tuning. (2) Prompt-tuning (Guo et al., 2022; Yang et al., 2023; Li et al., 2023; Dong et al., 2023), which involves creating prompts to address social biases.
@@ 3 Method
We propose a fine-grained bias mitigation method Fairness-Stamp (FAST). FAST operates through a two-step process, as depicted in Figure 3. In the first step, we propose to investigate if there are specific hidden states (i.e., layers) that play a more crucial role than others when recalling biased knowledge, as inspired by the knowledge localization works (Meng et al., 2022; Finlayson et al., 2021). Our biased knowledge localization is performed in three steps, biased run, counterfactual input and restoration run, with a complete description in Figure 4 in the Appendix B.1:
@@ 5 Conclusion
In this paper, we pioneer the fine-grained bias mitigation paradigm, which specifically focuses on human-relevant individual social biases/facts rather than broad group differences. We develop a novel evaluation benchmark BiasKE and propose the first Editable Fairness framework, FAST, capable of mitigating single social biases and scalable to mitigating thousands of biases concurrently. Extensive experiments across various models and datasets demonstrate the efficacy of our approach, showcasing its generalizability, specificity, and scalability. Our findings offer significant implications for future debiasing research. The limitation and future works can be referred to Appendix E.
@@ D.5 Computational Complexity Analysis
Appendix E Limitation and Future Works
@@
While our research yields important contributions, we acknowledge the presence of certain limitations. Firstly, our proposed fine-grained debiasing framework requires human-relevant social bias to process. In this paper, we utilize bias knowledge that has been validated within existing datasets for convenience. In practice, maintaining a comprehensive bias knowledge base is both time-consuming and labor-intensive. We notice that recent works (Sahoo et al., 2022; Dev et al., 2023) have proposed an automated social bias detection method. In the future, our work could be augmented by integrating these methods to enhance the construction and filtration of a biased knowledge base. Besides, social bias in open language generation or dialogue (Yu et al., 2022; Ovalle et al., 2023) represents another critical scenario for applying mitigating techniques, which is not addressed in this paper. Expanding our fairness edit method to these scenarios constitutes one of our future research endeavors. Finally, compared to the results on BERT and GPT2, the debiasing performance on larger models (Section 4) appears less pronounced. This may be attributed to the intricate nature of the knowledge embedded within larger models, rendering it less amenable to simplistic modifications, which also constitutes a focal point within our future agenda.

### v2

@@ 1 Introduction
Pre-trained Large Language Models (LLMs) have demonstrated exceptional performance on many tasks (Devlin et al., 2018; Floridi & Chiriatti, 2020; Brown et al., 2020). However, the encoded social stereotypes and human-like biases inevitably cause undesired behaviors when deploying LLMs in practice (Zhao et al., 2019; Navigli et al., 2023; Sheng et al., 2021). Existing approaches to mitigate biases in LLMs are mainly categorized into: (1) Fine-tuning (Zmigrod et al., 2019; Webster et al., 2020; He et al., 2022; Liang et al., 2020; Lauscher et al., 2021), which includes techniques such as re-balanced corpus pre-training, contrastive learning, projection methods, and efficient parameter tuning. (2) Prompt-tuning (Guo et al., 2022; Yang et al., 2023; Li et al., 2023b; Dong et al., 2023), which involves creating prompts to address social biases.
@@ 3 Method
We propose a fine-grained bias mitigation method Fairness-Stamp (FAST). FAST operates through a two-step process, as depicted in Figure 3. In the first step, we propose to investigate if there are specific hidden states (i.e., layers) that play a more crucial role than others when recalling biased knowledge, as inspired by the knowledge localization works (Meng et al., 2022a; Finlayson et al., 2021). Our biased knowledge localization is performed in three steps, biased run, counterfactual input and restoration run, with a complete description in Figure 4 in the Appendix B.1:
@@ 5 Conclusion
In this paper, we pioneer the fine-grained bias mitigation paradigm, which specifically focuses on human-relevant individual social biases/facts rather than broad group differences. We develop a novel evaluation benchmark BiasKE and propose the first Editable Fairness framework, FAST, capable of mitigating single social biases and scalable to mitigating thousands of biases concurrently. Extensive experiments across various models and datasets demonstrate the efficacy of our approach, showcasing its generalizability, specificity, and scalability. Our findings offer significant implications for future debiasing research. The limitation and future works can be referred to Appendix F.
@@ D.5 Computational Complexity Analysis
Appendix E Related Works
E.1 PLM Debiasing
Several approaches have been proposed for debiasing pre-trained language models. The techniques can be grouped into two categories. (1) Fine-tuning: This branch includes additional pre-training on re-balanced corpus Zmigrod et al. (2019); Webster et al. (2020) or with a contrastive objective He et al. (2022); Cheng et al. (2021), projection-based methods Liang et al. (2020); Ravfogel et al. (2020); Kaneko & Bollegala (2021); Dev et al. (2020) in the embedding space, in-training-based methods Han et al. (2021); He et al. (2022) and parameter-efficient fine-tuning Lauscher et al. (2021); Xie & Lukasiewicz (2023) methods. (2) Prompt-tuning: Prompt-tuning Guo et al. (2022); Yang et al. (2023); Li et al. (2023b); Dong et al. (2023) involves the generation of either discrete prompts or continuous prompts to mitigate social biases. There are also post-hoc approaches Schick et al. (2021) which are deployed in the inference phase. However, existing techniques treat social groups as interchangeable Gallegos et al. (2023). They seek to neutralize all protected attributes in the inputs or outputs of a model. These strategies tend to ignore or conceal distinct mechanisms of different social groups Hanna et al. (2020). In this paper, we develop evaluation and mitigation strategies that target specific historical biases, without defaulting to the erasure of social group identities as an adequate debiasing strategy.
E.2 Knowledge Locating
Localization aims to interpret a specific model component, including neurons, layers, or subnetworks Elhage et al. (2021); Rogers et al. (2021); Schneider & Vlachos (2021); Zeiler & Fergus (2014); Wang et al. (2022); Bolukbasi et al. (2021); Chen et al. (2024). For example, Dai et al. (2021) identifies a small set of knowledge neurons for each relational fact in BERT. Meng et al. (2022a) locate relational facts to middle FFN layers in autoregressive LLMs, specifically when the model processes the last token of the subject. In contemporaneous work, Bayazit et al. (2023) proposes a differentiable weight masking method to discover sparse subnetwork in GPT2 responsible for specific knowledge. Meng et al. (2022a) and Meng et al. (2022b) propose to utilize causal mediation analysis Vig et al. (2020); Finlayson et al. (2021) to identify individual layers and neurons that contribute to knowledge storing. Extending these ideas, we propose a two-step model-debiasing framework which firstly locates the key component responsible for storing specific biased knowledge and then calibrates the biased predictions.
E.3 Model Editing
Model Editing Sinitsin et al. (2020); De Cao et al. (2021) has been proposed to facilitate data-efficient modifications to model behavior while ensuring no detrimental impact on performance across other inputs. These approaches manipulate the model’s output for specific cases either by integrating external models with the original, unchanged model Mitchell et al. (2022); Murty et al. (2022); Dong et al. (2022); Hartvigsen et al. (2022); Huang et al. (2023); Zheng et al. (2023) or by altering the model parameters responsible for undesirable output Mitchell et al. (2021); Dai et al. (2021); Li et al. (2023a); Gupta et al. (2023); Hase et al. (2021); Meng et al. (2022a). The most relevant line of works in this regard is locate and edit Meng et al. (2022a; b); Dai et al. (2021); Li et al. (2023a), which suggests identifying neuron activations crucial to a model’s factual predictions and subsequently updating the feed-forward weights to edit the output. To the best of our knowledge, we are the first to use model editing techniques to achieve fine-grained model debiasing.
Appendix F Limitation and Future Works
@@
While our research yields important contributions, we acknowledge the presence of certain limitations. Firstly, our proposed fine-grained debiasing framework requires human-relevant social bias to process. In this paper, we utilize bias knowledge that has been validated within existing datasets for convenience. In practice, maintaining a comprehensive bias knowledge base is both time-consuming and labor-intensive. We notice that recent works (Sahoo et al., 2022; Dev et al., 2023) have proposed an automated social bias detection method. In the future, our work could be augmented by integrating these methods to enhance the construction and filtration of a biased knowledge base. Besides, expanding our fairness edit method against attack scenarios constitutes one of our future research endeavors Lyu et al. (2022; 2023). Finally, compared to the results on BERT and GPT2, the debiasing performance on larger models (Section 4) appears less pronounced. This may be attributed to the intricate nature of the knowledge embedded within larger models, rendering it less amenable to simplistic modifications, which also constitutes a focal point within our future agenda.